# Fractional attractors in light of the latest ACT observations

### v1

@@ Abstract
In light of the latest results from ACT observations we review a class of potentials labeled as fractional attractors, that can originate from Palatini gravity. We show in a model independent way that this class of potentials predicts both a spectral index $n_{s}$ and a tensor-to-scalar ratio $r$ which fit the $1\sigma$ region of the combination ACT+Planck data for a wide choice of the parameters. We also provide a numerical fit for the parameter space of this models in the case of a simple quadratic and quartic fractional potential.
@@ 1 Introduction
Among the others, the Palatini formulation of gravity [40, 41, 42] (and refs. therein) also provides a promising framework to understand inflation. In the Palatini formalism (e.g. considers the Levi-Civita connection a priori independent from the metric. When the gravity sector is non-minimally coupled to matter fields the phenomenological predictions differ from the prediction of metric gravity. In particular, the Palatini framework was shown to have many appealing features to build inflationary setups. An example is that of $F(R)$ models, for which the gravity sector is taken to be a general function of the Ricci scalar. This class of models generates asymptotically flat potentials that can be used to describe experimentally viable slow-roll inflation (e.g. [43, 44, 45] and refs. therein). A generalization of the $F(R)$ theories namely $F(R,X)$ can provide a class of fractional attractor potentials [46, 47], which exhibits an attractor behavior that generalizes the polynomial $\alpha$-attractors [31].
@@ 2 Fractional attractor potentials
with $U(\phi)$ an effective potential given by:
$U(\phi)=V(\phi)\left(1+8\alpha\frac{V(\phi)}{m^{4}_{P}}\right)^{-1}.$ (2)
This potential is a generalization of the polynomial $\alpha$-attractors studied in [31]. Such a setup can be built in the context of Palatini gravity [46] (see A).
@@
where $\phi_{end}$ is obtained by imposing the exit condition from slow-roll inflation [^1] The condition $|\eta|\simeq 1$ can also trigger the end of slow-roll. We checked that for the parameter space explored this condition never happens., i.e. $\epsilon(\phi)=1$. The field value $\phi_{N}$ at the time a given scale left the horizon is given by fixing $N$. From this, one can straightforwardly compute the CMB observables, i.e. the spectral index $n_{\textrm{s}}$ and the tensor-to-scalar ratio $r$:
@@
We can find a model independent expression for $r$ in the strong coupling limit $\alpha\to\infty$. In such a case, the potential asymptotically approaches the plateau $U(\phi)\rightarrow\frac{m^{4}_{P}}{8\alpha}$ for any specific choice of $V(\phi)$. Therefore, inserting this value in the numerator of Eq.(8) and using Eq.(7), it is straightforward to show that in a model-independent way
@@
This means that we can arbitrarily lower $r$ by increasing $\alpha$, in the same way of the quadratic Palatini models studied in [43]. In order to obtain a model-independent prediction for $n_{s}$ as well, we can consider a general potential $V(\phi)$. Any regular potential can be expanded in power series around $\phi=\phi_{N}$ as
$V(\phi)=\sum^{\infty}_{k=0}\frac{V^{(k)}(\phi_{N})}{k!}(\phi-\phi_{N})^{k},$ (10)
with $V^{(k)}(\phi_{N})$ the $k^{th}$ derivative of the potential evaluated at $\phi_{N}$. Hence, the leading contribution to any potential can be taken as a monomial around $\phi_{N}$. For this reason, we consider a potential in the form:
@@ Appendix A Palatini F(R,X)
This appendix is based on [46]. Consider the action

### v2

@@ Abstract
In light of the latest results from ACT observations we review a class of potentials labeled as fractional attractors, that can originate from Palatini gravity. We show that, for certain choices of the scalar potential $V(\phi)$, the fractional attractors predict both a spectral index $n_{s}$ and a tensor-to-scalar ratio $r$ that fall within the $1\sigma$ region of the combined ACT+Planck data for a wide range of parameters. We also provide a numerical fit for the parameter space of this models in the case of a simple quadratic and quartic fractional potential.
@@ 1 Introduction
Among the others, the Palatini formulation of gravity [40, 41, 42] (and refs. therein) also provides a promising framework to understand inflation. In the Palatini formalism the affine connection is a priori independent from the metric. When the gravity sector is non-minimally coupled to matter fields the phenomenological predictions differ from the prediction of metric gravity. In particular, the Palatini framework was shown to have many appealing features to build inflationary setups. An example is that of $F(R)$ models, for which the gravity sector is taken to be a general function of the Ricci scalar. This class of models generates asymptotically flat potentials that can be used to describe experimentally viable slow-roll inflation (e.g. [43, 44, 45] and refs. therein). A generalization of the $F(R)$ theories namely $F(R,X)$ can provide a class of fractional attractor potentials [46, 47], which exhibits an attractor behavior that generalizes the polynomial $\alpha$-attractors [31].
@@ 2 Fractional attractor potentials
with $U(\phi)$ the inflaton potential. It is well known that data [6, 9] strongly favors nearly flat inflaton potentials. One of the simplest way to construct an asymptotically flat potential is via the parametrization
$U(\phi)=V(\phi)\left(1+8\alpha\frac{V(\phi)}{m^{4}_{P}}\right)^{-1},$ (2)
where $\alpha$ is a dimensionless parameter and $V(\phi)$ a generating function with the dimensions of a scalar potential. When $\alpha\to\infty$, the potential $U$ becomes asymptotically flat with $U(\phi)\approx\frac{m^{4}_{P}}{8\alpha}$ for any specific choice of $V(\phi)$. The potential in eq. (2) is a generalization of the polynomial $\alpha$-attractors studied in [31]. Such a setup can be built in the context of Palatini gravity [46] (see A).
@@
where $\phi_{end}$ is obtained by imposing the exit condition from slow-roll inflation [^1] The condition $|\eta|\simeq 1$ can also trigger the end of the slow-roll approximation. We explicitly checked that during the inflationary evolution, i.e. until $\epsilon=1$ , the condition $|\eta|\ll 1$ was satisfied for the whole parameter space considered., i.e. $\epsilon(\phi)=1$. The field value $\phi_{N}$ at the time a given scale left the horizon is given by fixing $N$. From this, one can straightforwardly compute the CMB observables, i.e. the spectral index $n_{\textrm{s}}$ and the tensor-to-scalar ratio $r$:
@@
We can find a model independent expression for $r$ in the strong coupling limit $\alpha\to\infty$. As mentioned before, the potential asymptotically approaches the plateau $U(\phi)\rightarrow\frac{m^{4}_{P}}{8\alpha}$ for any specific choice of $V(\phi)$. Therefore, inserting this value in the numerator of Eq.(8) and using Eq.(7), it is straightforward to show that in a model-independent way
@@
This means that we can arbitrarily lower $r$ by increasing $\alpha$, in the same way of the quadratic Palatini models studied in [43].
On the other hand, the predictions for $n_{s}$ strongly depend on the choice of $V(\phi)$. As reference cases, we consider potentials in the following monomial form:
@@ Appendix A Palatini F(R,X)
In this appendix we briefly describe a possible setup to generate a potential in the form of fractional attractors [^2] This setup was first introduced in [46] as a generalization of the Palatini $F(R)$ models, in order to solve inflationary issues beyond slow-roll for models containing $R^{n}$ terms with $n>2$.. Consider the action